# Single vibronic level fluorescence spectra from Hagedorn wavepacket dynamics

### v2

@@ III.1 Two-dimensional harmonic potentials
is a rotation matrix. The analysis of spectra here becomes more challenging, even in two-dimensional systems. We observe significant differences in the intensity pattern compared to unrotated systems. In our particular case, the frequency range of the spectral region also appears to be decreased. From a time-independent perspective, this could be explained by the fact that the Duschinsky rotation reduces the Franck-Condon overlaps between the initial wavepacket and higher vibrational states on the ground electronic surface. It is important to note that in this most general harmonic model, the Hagedorn wavepackets do not maintain the separable form of (10) during propagation, since $Q_{t}$ is no longer diagonal when evolved under a Duschinsky-rotated potential. With Hagedorn’s ingenious construction, the propagation algorithm maintains its simplicity even for general, non-separable Hagedorn wavepackets, which are not simply tensor products of eigenfunctions of one-dimensional harmonic oscillators.
All spectra shown in Fig. 2 were obtained with the Hagedorn approach and are validated against quantum split operator results in Fig. 3. In contrast to the Hagedorn approach, the quantum dynamics must be performed separately for each initial vibrational level in each system. Where applicable, the Hagedorn results are also compared to the spectra evaluated using the TGA[36, 68, 38] (for $1^{0}2^{0}$) and its extended version[44, 22] (for $1^{1}2^{0}$ and $1^{1}2^{1}$). Before calculating the differences, the spectra were all scaled and shifted as in Fig. 2. As expected, we observe excellent agreement between the methods. The comparison with the (extended) TGA results serves as a further validation, as these calculations are performed in Heller’s Gaussian parametrization[36, 69] that differs substantially from Hagedorn’s. In addition, ETGA employs a different approach to propagate the linear polynomial prefactor of the Gaussian.
@@
In higher dimensions, the quantum split-operator approach quickly becomes unfeasible due to the curse of dimensionality, particularly when treating excited vibrational wavefunctions for which denser grids are generally needed. In contrast, Hagedorn wavepackets circumvent the need for a grid and can be efficiently propagated within a harmonic approximation using the same trajectory as the simple thawed Gaussian. With the algebraic expressions developed in Ref. 67, we can now use Hagedorn wavepackets in much higher dimensions.
Figure 4 shows SVL emission spectra in a 100-dimensional system that incorporates displacement, distortion, and Duschinsky rotation effects (all parameters are specified in the supplementary material). Except for the potential energy surface and initial wavepackets, the computation parameters (time step, total time of propagation, frequency of evaluation of the autocorrelation function, broadening of the spectra) are identical to those for the two-dimensional systems. We choose two modes (namely modes 30 and 40) to be vibrationally excited in the initial wavepackets, computing emission spectra from initial vibrational levels $30^{a}40^{b}$, with the same set of $(a,b)$ as in the two-dimensional cases.
@@
The spectra become significantly more complex in higher-dimensional systems. In this particular example, similarity can be observed between the $30^{0}40^{0}$ and $30^{1}40^{0}$ spectra as well as between the $30^{1}40^{1}$ and $30^{2}40^{0}$ spectra. However, significant changes in the intensity pattern can be seen when comparing spectra with different excitations in mode 40 (compare $30^{0}40^{0}$ and $30^{0}40^{3}$, or $30^{1}40^{0}$ and $30^{1}40^{1}$), indicating that the excitation in mode 40 likely has a more pronounced impact on the overall spectrum compared to the excitation in mode 30. We also observe broadened tails in the lower frequency region, which are likely attributable to the effects of Duschinsky rotation here (compare to Fig. 5 where the Duschinsky couplings are neglected). As a time-dependent approach, Hagedorn wavepacket dynamics can intrinsically take into account Duschinsky rotation effects, in contrast to the time-independent method, which faces challenges related to the pre-selection of likely transitions and the computation of non-separable Franck-Condon integrals in high-dimensional systems.
Figure 5: SVL emission spectra from different vibrational levels (indicated in each panel) in a 100-dimensional two-state displaced and distorted harmonic system (same as in Fig. 4 but without Duschinsky rotation).
@@ IV Conclusions
In conclusion, Hagedorn wavepackets are well suited for simulating SVL emission spectra. While the present work only considered model potentials, Hagedorn wavepackets can readily be applied to realistic molecular systems using harmonic potentials constructed from ab initio electronic structure calculations. Moreover, as in the thawed Gaussian approximation for computing ground-state emission spectra, anharmonicity effects could be partially captured by propagating Hagedorn wavepackets within local harmonic approximation using on-the-fly ab initio data.
@@ Acknowledgements.
The authors acknowledge the financial support from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 683069–MOLEQULE) and from the EPFL.

### v3

@@ III.1 Two-dimensional harmonic potentials
is a rotation matrix. The analysis of spectra here becomes more challenging, even in two-dimensional systems. Significant differences in the intensity pattern appear compared to unrotated systems. In our particular case, the frequency range of the spectral region also decreases. From a time-independent perspective, this could be explained by the fact that the Duschinsky rotation reduces the Franck-Condon overlaps between the initial wavepacket and higher vibrational states on the ground electronic surface. Finally, in this most general harmonic model, the propagated Hagedorn wavepackets do not retain the simple, separable form of products of one-dimensional Hermite functions as in Eq. (11), since $Q_{t}$ is no longer diagonal when evolved under a Duschinsky-rotated potential. With Hagedorn’s ingenious construction, the propagation algorithm maintains its simplicity even for general, non-separable Hagedorn wavepackets, which are not simple products of eigenfunctions of one-dimensional harmonic oscillators.
All spectra shown in Fig. 2 were obtained with the Hagedorn approach and are validated against quantum split operator results in Fig. 3. In contrast to the Hagedorn approach, the quantum dynamics must be performed separately for each initial vibrational level in each system. Where applicable, the Hagedorn results are also compared to the spectra evaluated using the TGA[36, 68, 38] (for $1^{0}2^{0}$) and its extended version[44, 22] (for $1^{1}2^{0}$ and $1^{1}2^{1}$). Before calculating the differences, the spectra were all scaled and shifted as in Fig. 2. As expected, we observe excellent agreement between the methods. The comparison with the (extended) TGA results serves as a further validation, as these calculations were performed in Heller’s Gaussian parametrization[36, 69] that differs substantially from Hagedorn’s. In addition, ETGA employs a different approach to propagate the linear polynomial prefactor of the Gaussian.
@@
In higher dimensions, the quantum split-operator approach quickly becomes unfeasible due to the curse of dimensionality, particularly when treating excited vibrational wavefunctions for which denser grids are generally needed. In contrast, Hagedorn wavepackets circumvent the need for a grid and can be efficiently propagated within a harmonic approximation using the same trajectory as the guiding thawed Gaussian. With the algebraic expressions from Ref. 67, we can now use Hagedorn wavepackets in much higher dimensions.
Figure 4 shows SVL emission spectra in a 100-dimensional system that incorporates displacement, distortion, and Duschinsky rotation effects (all parameters are specified in the supplementary material). The time step, total time of propagation, frequency of evaluation of the autocorrelation function, and broadening of the spectra are identical to those for the two-dimensional systems. We choose two modes (namely modes 31 and 41) to be vibrationally excited in the initial wavepackets, computing emission spectra from initial vibrational levels $31^{a}41^{b}$, with the same set of $(a,b)$ as in the two-dimensional cases.
@@
The spectra become significantly more complex in higher-dimensional systems. In this example, similarity is observed between the $31^{0}41^{0}$ and $31^{1}41^{0}$ spectra as well as between the $31^{1}41^{1}$ and $31^{2}41^{1}$ spectra. However, significant changes in the intensity pattern appear when comparing spectra with different excitations in mode 41 (compare $31^{0}41^{0}$ and $31^{0}41^{3}$, or $31^{1}41^{0}$ and $31^{1}41^{1}$), suggesting that the excitation in mode 41 has a more pronounced impact on the overall spectrum compared to the excitation in mode 31. We also observe broadened features in the lower frequency region, which are likely attributable to the effects of Duschinsky rotation here (compare to Fig. S1 where the Duschinsky couplings are neglected).
The initial vibrational excitation in SVL spectra can expand the spectral range and induce transitions to a large number of highly excited vibrational states of the ground electronic state. Considering, among many other possibilities, only the final vibrational states with five vibrational quanta distributed in three vibrational modes already requires evaluating $970200$ transitions (see supplementary material for derivation) at wavenumbers $>-20000\,\text{cm}^{-1}$, compared to $10000$ evaluations of the autocorrelation function used to generate the spectra in Fig. 4. The actual number of possible transitions is much higher, given the possibility of more quanta distributed among more modes. As a consequence, a sophisticated prescreening is needed. Being a time-dependent approach, Hagedorn wavepacket dynamics can directly capture the broadened feature without the need to compute the huge number of individual transitions that contribute to the tail. In contrast, the time-independent approach can identify contributions from specific transitions, but this is less critical at the relatively low spectral resolution in which we are interested.
[…]
@@ IV Conclusions
In conclusion, Hagedorn wavepackets are well suited for simulating SVL emission spectra. While the present work only considered model potentials, Hagedorn wavepackets can be readily applied to realistic molecular systems using harmonic potentials constructed from ab initio electronic structure calculations. Moreover, as in the thawed Gaussian approximation for computing ground-state emission spectra, anharmonicity effects could be partially captured by propagating Hagedorn wavepackets within local harmonic approximation using on-the-fly ab initio data.
@@
Supplementary material
The supplementary material[89] contains the spectra in the 100-dimensional system without Duschinsky couplings, the parameters of the harmonic systems analyzed in Section III.2, the comparison with the time-independent approach for the $0\,\text{cm}^{-1}$ peaks in Fig. 4, as well as the Python program used to compute the overlaps between Hagedorn functions.[67]
@@ Acknowledgements.
The authors acknowledge the financial support from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 683069–MOLEQULE).